OpenStudy (anonymous):

Is this an identity or no solution? 6-6x=-6x-6 Original Problem: 3(2-2x)=-6(x-1)

OpenStudy (usukidoll):

6-6x = -6x+6 it's an identity just the right side is written differently

jimthompson5910 (jim_thompson5910):

6-6x=-6x-6 6-6x+6x=-6x-6+6x ... add 6x to both sides 6 + 0x = 0x - 6 6 + 0 = 0 - 6 6 = -6 ... which is always false so there are no solutions (this is a contradiction)

OpenStudy (anonymous):

So is it no solution?

jimthompson5910 (jim_thompson5910):

however, for the original equation 3(2-2x)=-6(x-1) 3(2)+3(-2x)=-6(x)-6(-1) ... distribute (remember to distribute to EVERY term inside) 6 - 6x = -6x + 6 6 - 6x = 6 - 6x and we can see the two sides are identical...so we have an identity which means we have infinitely many solutions

OpenStudy (usukidoll):

UGH YOU SCARED ME MAN! DON'T DO THAT >:O

jimthompson5910 (jim_thompson5910):

Notice how 6-6x=-6x-6 (what you posted first) and 6 - 6x = -6x + 6 (what you get when you distribute) are different

jimthompson5910 (jim_thompson5910):

no, 6-6x=-6x-6 is different from 6 - 6x = -6x + 6

OpenStudy (usukidoll):

so it is no solution.. my bad ^^

OpenStudy (usukidoll):

the negative sign on the right threw everything off when you distribute

jimthompson5910 (jim_thompson5910):

If the original problem is 3(2-2x)=-6(x-1), then you distribute things out to get 6 - 6x = -6x + 6

jimthompson5910 (jim_thompson5910):

That is an identity which leads you to infinitely many solutions (basically all real numbers) but if the original is 6-6x=-6x-6, then it's going to lead to a contradiction